The decomposition of a certain mass of CaCO3 gave 11.2 dm3 of CO2 gas at STP. The mass of KOH required to completely neutralise the gas is
(A) 56 g
(B) 28 g
(C) 42 g
(D) 20 g

Grade:12

2 Answers

R Shyam Sundar
4 years ago
For complete neutralisation,
2KOH+CO2 gives K2CO3 + H2O
For 11.2dm^3 i.e.,11.2L(half mole) we need 1mole KOH.
So weight of KOH is 56g
 
Vikas TU
4 years ago
Dear student 
Weight of 11.2 dm^3 of CO2 gas at STP = 44/2 = 22 gram 
KOH + CO2 -> KHCO3 
KOH required = 56/44  * 22 = 28 gram 
B is correct ans.
Hope this helps 
Good Luck